Grade 12th passMechanics

Work done in lifiting up a stone of mass 2kg to 4metres is

Sushant sharma
8 Years agoGrade 12th pass

1 Answer

Arun
8 Years ago
Dear student
Force = m* g Newton
F = 2 * 9.8 = 19.6 N
Now
Work done = m*g*h
= 2* 9.8 * 4 = 78.4 Joules
Regards
Arun (askIITians forum expert)